OpenStudy (anonymous):

If g(x)=x^2-2x-1,find g(-3) a. –22 c. 14 b. 2 d. 16

hartnn (hartnn):

to get g(-3) just plug in x=-3 in g(x) !

hartnn (hartnn):

\( (-3)^2-2(-3)-1 = ...? \)

OpenStudy (anonymous):

14

hartnn (hartnn):

yes, thats correct :)

OpenStudy (anonymous):

Which of the following is a horizontal stretch of the parent graph? f(x)=x a.g(x)=5x b.g(x)=5x c.g(x)=1/5x d.g(x)=1/5x